OpenStudy (anonymous):

if w=6 when z=2 find w when z=-3

OpenStudy (anonymous):

6 : 2 x : -3 ratios. cross multiply

OpenStudy (anonymous):

since 6 = 3(2) we can see that the equation to find w is w=3z. using this information we can plug in -3 and see that: w=3(-3) w=-9

OpenStudy (anonymous):

So the answer is +9

OpenStudy (anonymous):

why is it +9?

OpenStudy (anonymous):

It is the answer on that was listed on the computer from the teacher. if we set it up with basic ratio and multiply -6*2 and x*3 = 2x/2=18/2 x=9

OpenStudy (anonymous):

when you cross multiply, you get 6*-3/2 = -18/2 = -9, not +9

OpenStudy (anonymous):

Hmm, interesting. The teacher only had 9 for the answer.

OpenStudy (anonymous):

your teacher is wrong

OpenStudy (anonymous):

logic follows logic, not what your teacher says

OpenStudy (anonymous):

as long as you follow the correct logic, you will always end up with the correct answer.